A resistor develops heat at the rate of 20W when the potential difference across its ends is 30V. What is the resistance of the resistor?


1
Expert's answer
2021-02-08T18:50:58-0500

"P=\frac{V^2}{R}"


"P=20\:W"


"V=30v"


"20=\frac{30^2}{R}= \frac{900}{R}"


"R=\frac{900}{20}=45\Omega"